a 1200-kg car goes from 10m/s to 24 m/s in 16s. What is the average force acting upon it
Answer by fractalier(6550)

F = ma = m * (change in velocity / change in time)
F = (1200) * ((24-10) / 16)
F = 1200 * 14/16 =
F = 1050 N